Personal tools
Document Actions

Work Log

Latest Data Historical Data Station Info Station Photos Work Log
Initials Who
SH Sara Haines
CBC Chris Calloway
 
Station Events and Changes
Date/Time Who Description
5/20/2014 SH Redeployed after manufacturer repairs.
3/11/2014 SH SODAR was damaged during the snow storm of Jan/Feb 2014. It has been taken offline until repairs can be made.
8/28/2013 CBC Stopped measurement at 14:20 UTC to demonstrate settings to Megan Schutt. Did no turn back on for 12 hours.
7/9/2013 CBC Changed beam angle back to narrow and max reflection height to 200m.
7/2/2013 CBC Changed to wide beam angle.
6/18/2013 CBC Changed height of maximum reflections to wait for to 0m.
5/13/2013 CBC Trimmed the legs off the antenna stand. The lip of the antenna enclosure is now even with the fencing. Conducted a power consumption survey.
3/12/2013 CBC Change beam tilt back to narrow beam at 14:30 UTC (note DST shift since last config change).
2/26/2013 CBC Change configuration from narrow to wide beam.
2/7/2013 CBC Plot generation restored.
2/6/2013 CBC NetCDF processing restored.
2/5/2013 CBC Raw data transfer to repository restored.
2/1/2013 CBC Enabled RDP and SFTP on new sodar control laptop. Reployment complete.
1/31/2013 CBC Replaced Antenna to SPU cables. Device properly functioning.
1/30/2013 CBC Sodar SPU not powering up. Consult Scintec.
1/29/2013 CBC Assembled new sodar platform at Billy Mitchell
1/28/2013 CBC Commence redeployment field work. Built new sodar platform at CSI.
1/27/2013 CBC Hwy 12 access to sodar site restored.
1/22/2013 CBC Replacement sodar control laptop received. Applied OS (Win 7) updates. Installed API Run 1.43.
1/16/2013 CBC Ordered replacement sodar control laptop.
1/14/2013 CBC Receive research permit renewal from NPS.
1/8/2013 CBC Isolated laptop failure to receive section of UART. Recommend repair or replacement.
12/21/2013 CBC Submit annual investigator's report to NPS and apply for research permit renewal.
10/29/2013 CBC Physical access to sodar site blocked by storm damage to Hwy 12.
10/2/2013 CBC Isolate communications failure to sodar control laptop in lab. Required boring out one retaining screw on SPU lid to test SPU internally.
9/26/2012 CBC Removed sodar equipment from Billy Mitchell to Seim lab.
7/26/2012 CBC Stopped at restarted the device server twice starting at 13:36:29 UTC. SPU will not connect. Notified Scintec of error id codes i350 and c175.
7/23/2012 CBC SPU communications errors started at 12:06:15 UTC. SPU not connected errors start filling logs at one minute intervals.
6/15/2012 CBC Partially restored NetCDF and plot generation by creating new data processing profile starting on 4/16/2012 because base observation altitude and number of bins changed with new configuration.
5/16/2012 CBC Took measurement offline at 14:53 UTC. Turned power back to max for all measurements. Measurement running again at 14:58 UTC.
5/2/2012 CBC Took measurement offline at 13:50 UTC. Turned on low power (32% volume) for all measurements for testing. Restarted measurement at 13:52 UTC.
4/16/2012 CBC Took measurement offline at 18:46 UTC. Ran self-test (passed). Reset all settings. Selected narrow angle and optimized pulse range (both recommended). Set wait for echoes to 0m for both horizontal and vertical subcycles. This change effectively resets pulse durations to defaults. Saved settings. Restarted measurement at 19:50 UTC.
4/11/2012 CBC Took measurement offline at 14:09 UTC to examine primary settings. Restarted measurement at 14:19 UTC.
12/12/2011 CBC Took measurement offline at 21:03 UTC. Kept previous settings and adding waiting for echoes from maximum height between emissions. Restarted measurement at 21:08 UTC.
10/28/2011 CBC Took measurement offline at 23:56 UTC. Kept previous settings and changed duration setting on all emission frequencies and all subcycles to 20m. Restarted measurement at 00:00 UTC.
10/19/2011 CBC Took measurement offline at 19:36 UTC. Reset all settings and changed beams to narrow angle. Restarted measurement at 19:38 UTC.
10/14/2011 CBC Stopped measurement from 19:41 through 19:46 UTC to demonstrate new settings to HS.
10/13/2011 CBC Consulted with Danial Layfield at Scintec. Took measurement offline at 20:00 UTC. Changed duration setting on all emission frequencies and all subcycles to 20m and max height setting back to 200m. Restarted measurement at 21:05 UTC.

Edit this table


Powered by Plone, the Open Source Content Management System